\begin{thebibliography}{37}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Achiam et~al.(2023)Achiam, Adler, Agarwal, Ahmad, Akkaya, Aleman,
  Almeida, Altenschmidt, Altman, Anadkat, et~al.]{gpt4}
J.~Achiam, S.~Adler, S.~Agarwal, L.~Ahmad, I.~Akkaya, F.~L. Aleman, D.~Almeida,
  J.~Altenschmidt, S.~Altman, S.~Anadkat, et~al.
\newblock {GPT}-4 technical report.
\newblock \emph{arXiv preprint arXiv:2303.08774}, 2023.

\bibitem[Anthropic(2024)]{anthropic2024claude3}
Anthropic.
\newblock The claude 3 model family: Opus, sonnet, haiku, 2024.

\bibitem[Arora and Barak(2009)]{arora2009computational}
S.~Arora and B.~Barak.
\newblock \emph{Computational complexity: a modern approach}.
\newblock Cambridge University Press, 2009.

\bibitem[Bhattamishra et~al.(2020)Bhattamishra, Patel, and Goyal]{bhat}
S.~Bhattamishra, A.~Patel, and N.~Goyal.
\newblock On the computational power of transformers and its implications in
  sequence modeling.
\newblock In \emph{Proceedings of the 24th Conference on Computational Natural
  Language Learning}, pages 455--475, 2020.

\bibitem[Davis et~al.(1994)Davis, Sigal, and Weyuker]{davis1994computability}
M.~Davis, R.~Sigal, and E.~J. Weyuker.
\newblock \emph{Computability, complexity, and languages: fundamentals of
  theoretical computer science}.
\newblock Elsevier, 1994.

\bibitem[Dubey et~al.(2024)Dubey, Jauhri, Pandey, Kadian, Al-Dahle, Letman,
  Mathur, Schelten, Yang, Fan, et~al.]{dubey2024llama}
A.~Dubey, A.~Jauhri, A.~Pandey, A.~Kadian, A.~Al-Dahle, A.~Letman, A.~Mathur,
  A.~Schelten, A.~Yang, A.~Fan, et~al.
\newblock The llama 3 herd of models.
\newblock \emph{arXiv preprint arXiv:2407.21783}, 2024.

\bibitem[Feng et~al.(2023)Feng, Zhang, Gu, Ye, He, and Wang]{feng}
G.~Feng, B.~Zhang, Y.~Gu, H.~Ye, D.~He, and L.~Wang.
\newblock Towards revealing the mystery behind chain of thought: a theoretical
  perspective.
\newblock \emph{Advances in Neural Information Processing Systems},
  36:\penalty0 70757--70798, 2023.

\bibitem[Gemini et~al.(2023)Gemini, Anil, Borgeaud, Alayrac, Yu, Soricut,
  Schalkwyk, Dai, Hauth, Millican, et~al.]{team2023gemini}
T.~Gemini, R.~Anil, S.~Borgeaud, J.-B. Alayrac, J.~Yu, R.~Soricut,
  J.~Schalkwyk, A.~M. Dai, A.~Hauth, K.~Millican, et~al.
\newblock Gemini: a family of highly capable multimodal models.
\newblock \emph{arXiv preprint arXiv:2312.11805}, 2023.

\bibitem[Goldblum et~al.(2024)Goldblum, Finzi, Rowan, and
  Wilson]{GoldblumFRW24}
M.~Goldblum, M.~A. Finzi, K.~Rowan, and A.~G. Wilson.
\newblock Position: The no free lunch theorem, kolmogorov complexity, and the
  role of inductive biases in machine learning.
\newblock 2024.

\bibitem[Guo et~al.(2025)Guo, Yang, Zhang, Song, Zhang, Xu, Zhu, Ma, Wang, Bi,
  et~al.]{deepseek}
D.~Guo, D.~Yang, H.~Zhang, J.~Song, R.~Zhang, R.~Xu, Q.~Zhu, S.~Ma, P.~Wang,
  X.~Bi, et~al.
\newblock Deepseek-r1: Incentivizing reasoning capability in llms via
  reinforcement learning.
\newblock \emph{arXiv preprint arXiv:2501.12948}, 2025.

\bibitem[Hearn and Demaine(2005)]{hearn2005pspace}
R.~A. Hearn and E.~D. Demaine.
\newblock Pspace-completeness of sliding-block puzzles and other problems
  through the nondeterministic constraint logic model of computation.
\newblock \emph{Theoretical Computer Science}, 343\penalty0 (1-2):\penalty0
  72--96, 2005.

\bibitem[Hendrycks et~al.(2021)Hendrycks, Burns, Kadavath, Arora, Basart, Tang,
  Song, and Steinhardt]{hendrycks2021measuring}
D.~Hendrycks, C.~Burns, S.~Kadavath, A.~Arora, S.~Basart, E.~Tang, D.~Song, and
  J.~Steinhardt.
\newblock Measuring mathematical problem solving with the math dataset.
\newblock \emph{arXiv preprint arXiv:2103.03874}, 2021.

\bibitem[Hutter(2005)]{uai}
M.~Hutter.
\newblock \emph{Universal artificial intelligence: Sequential decisions based
  on algorithmic probability}.
\newblock Springer Science \& Business Media, 2005.

\bibitem[Impagliazzo and Paturi(2001)]{impagliazzo2001complexity}
R.~Impagliazzo and R.~Paturi.
\newblock On the complexity of k-sat.
\newblock \emph{Journal of Computer and System Sciences}, 62\penalty0
  (2):\penalty0 367--375, 2001.

\bibitem[Jimenez et~al.(2023)Jimenez, Yang, Wettig, Yao, Pei, Press, and
  Narasimhan]{jimenez2023swe}
C.~E. Jimenez, J.~Yang, A.~Wettig, S.~Yao, K.~Pei, O.~Press, and K.~Narasimhan.
\newblock Swe-bench: Can language models resolve real-world github issues?
\newblock \emph{arXiv preprint arXiv:2310.06770}, 2023.

\bibitem[Lecun(2024)]{lecun}
Y.~Lecun.
\newblock Meta ai, open source, limits of llms, agi \& the future of ai, 2024.
\newblock \url{https://www.youtube.com/watch?v=5t1vTLU7s40}.

\bibitem[Levin(1973)]{levin1}
L.~A. Levin.
\newblock Universal sequential search problems.
\newblock \emph{Problemy peredachi informatsii}, 9\penalty0 (3):\penalty0
  115--116, 1973.

\bibitem[Levin(1984)]{levin2}
L.~A. Levin.
\newblock Randomness conservation inequalities; information and independence in
  mathematical theories.
\newblock \emph{Information and Control}, 61\penalty0 (1):\penalty0 15--37,
  1984.

\bibitem[Li et~al.(2024)Li, Liu, Zhou, and Ma]{tengyu}
Z.~Li, H.~Liu, D.~Zhou, and T.~Ma.
\newblock Chain of thought empowers transformers to solve inherently serial
  problems.
\newblock In \emph{The Twelfth International Conference on Learning
  Representations, {ICLR} 2024, Vienna, Austria, May 7-11, 2024}, 2024.

\bibitem[Merrill and Sabharwal(2023)]{ashish2}
W.~Merrill and A.~Sabharwal.
\newblock The parallelism tradeoff: Limitations of log-precision transformers.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  11:\penalty0 531--545, 2023.

\bibitem[Merrill and Sabharwal(2024)]{ashish}
W.~Merrill and A.~Sabharwal.
\newblock The expressive power of transformers with chain of thought.
\newblock In \emph{The Twelfth International Conference on Learning
  Representations, {ICLR} 2024, Vienna, Austria, May 7-11, 2024}, 2024.

\bibitem[Merrill and Sabharwal(2025)]{depth}
W.~Merrill and A.~Sabharwal.
\newblock A little depth goes a long way: The expressive power of log-depth
  transformers.
\newblock \emph{CoRR}, abs/2503.03961, 2025.

\bibitem[P{\'{e}}rez et~al.(2021)P{\'{e}}rez, Barcel{\'{o}}, and
  Marinkovic]{perez}
J.~P{\'{e}}rez, P.~Barcel{\'{o}}, and J.~Marinkovic.
\newblock Attention is turing-complete.
\newblock \emph{J. Mach. Learn. Res.}, 22:\penalty0 75:1--75:35, 2021.
\newblock URL \url{https://jmlr.org/papers/v22/20-302.html}.

\bibitem[Pettorossi(2014)]{pettorossi2014elements}
A.~Pettorossi.
\newblock \emph{Elements of computability, decidability, and complexity}.
\newblock Aracne editrice Srl, 2014.

\bibitem[Post(1936)]{post1}
E.~L. Post.
\newblock Finite combinatory processes—formulation1.
\newblock \emph{The journal of symbolic logic}, 1\penalty0 (3):\penalty0
  103--105, 1936.

\bibitem[Qiu et~al.(2024)Qiu, Xu, Bao, and Tong]{prompt}
R.~Qiu, Z.~Xu, W.~Bao, and H.~Tong.
\newblock Ask, and it shall be given: Turing completeness of prompting.
\newblock \emph{CoRR}, abs/2411.01992, 2024.

\bibitem[Radford et~al.(2019)Radford, Wu, Child, Luan, Amodei, Sutskever,
  et~al.]{radford2019language}
A.~Radford, J.~Wu, R.~Child, D.~Luan, D.~Amodei, I.~Sutskever, et~al.
\newblock Language models are unsupervised multitask learners.
\newblock \emph{OpenAI blog}, 1\penalty0 (8):\penalty0 9, 2019.

\bibitem[Rein et~al.(2023)Rein, Hou, Stickland, Petty, Pang, Dirani, Michael,
  and Bowman]{rein2023gpqa}
D.~Rein, B.~L. Hou, A.~C. Stickland, J.~Petty, R.~Y. Pang, J.~Dirani,
  J.~Michael, and S.~R. Bowman.
\newblock Gpqa: A graduate-level google-proof q\&a benchmark.
\newblock \emph{arXiv preprint arXiv:2311.12022}, 2023.

\bibitem[Schmidhuber(2004)]{oops}
J.~Schmidhuber.
\newblock Optimal ordered problem solver.
\newblock \emph{Machine Learning}, 54:\penalty0 211--254, 2004.

\bibitem[Schuurmans et~al.(2024)Schuurmans, Dai, and
  Zanini]{schuurmans2024autoregressive}
D.~Schuurmans, H.~Dai, and F.~Zanini.
\newblock Autoregressive large language models are computationally universal.
\newblock \emph{arXiv preprint arXiv:2410.03170}, 2024.

\bibitem[Sistla and Clarke(1985)]{sistla1985complexity}
A.~P. Sistla and E.~M. Clarke.
\newblock The complexity of propositional linear temporal logics.
\newblock \emph{Journal of the ACM (JACM)}, 32\penalty0 (3):\penalty0 733--749,
  1985.

\bibitem[Strobl et~al.(2024)Strobl, Merrill, Weiss, Chiang, and
  Angluin]{strobl2024formal}
L.~Strobl, W.~Merrill, G.~Weiss, D.~Chiang, and D.~Angluin.
\newblock What formal languages can transformers express? a survey.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  12:\penalty0 543--561, 2024.

\bibitem[Umans(2025)]{pmtm2}
C.~Umans.
\newblock Decidability and tractability: Solution set 3, 2025.
\newblock \url{https://users.cms.caltech.edu/~umans/cs21/soln3.pdf}.

\bibitem[Williams(2025)]{ryan}
R.~R. Williams.
\newblock Simulating time with square-root space.
\newblock \emph{arXiv preprint arXiv:2502.17779}, 2025.

\bibitem[Yang et~al.(2024)Yang, Chiang, and Angluin]{DBLP:conf/nips/Yang0A24}
A.~Yang, D.~Chiang, and D.~Angluin.
\newblock Masked hard-attention transformers recognize exactly the star-free
  languages.
\newblock In A.~Globersons, L.~Mackey, D.~Belgrave, A.~Fan, U.~Paquet, J.~M.
  Tomczak, and C.~Zhang, editors, \emph{Advances in Neural Information
  Processing Systems 38: Annual Conference on Neural Information Processing
  Systems 2024, NeurIPS 2024, Vancouver, BC, Canada, December 10 - 15, 2024},
  2024.

\bibitem[Yang et~al.(2025)Yang, Srebro, McAllester, and Li]{pencil}
C.~Yang, N.~Srebro, D.~McAllester, and Z.~Li.
\newblock Pencil: Long thoughts with short memory.
\newblock \emph{arXiv preprint arXiv:2503.14337}, 2025.

\bibitem[Zubi{\'c} et~al.(2024)Zubi{\'c}, Sold{\'a}, Sulser, and
  Scaramuzza]{zubic}
N.~Zubi{\'c}, F.~Sold{\'a}, A.~Sulser, and D.~Scaramuzza.
\newblock Limits of deep learning: Sequence modeling through the lens of
  complexity theory.
\newblock \emph{arXiv preprint arXiv:2405.16674}, 2024.

\end{thebibliography}

\documentclass[11pt]{article}

\usepackage{float}
\usepackage{array, makecell} %
\usepackage{xcolor,colortbl}
\usepackage{framed,color}
\usepackage{xcolor}
\usepackage{enumitem}
\usepackage{thm-restate}
\usepackage{natbib}
\usepackage{booktabs}       

\usepackage{amsthm,amsmath,amssymb}
\usepackage[bookmarks=true,hypertexnames=false,pagebackref]{hyperref}
\hypersetup{colorlinks=true, citecolor=blue, linkcolor=red,
  urlcolor=blue}
\usepackage{newpxtext}
\usepackage[libertine,vvarbb]{newtxmath}

\usepackage[T1]{fontenc} 
\usepackage{textcomp} 

\usepackage[scr=rsfso]{mathalfa}
\usepackage{nicefrac}
\usepackage{cleveref}
\usepackage{graphicx}
\usepackage{aliascnt}
\usepackage[section]{placeins}
\usepackage{tcolorbox}
\usepackage{epigraph}
\usepackage{xifthen}
\usepackage{latexsym}
\usepackage{framed}
\usepackage{fullpage}
\usepackage{bm}
\usepackage{braket}

\newtheorem{theorem}{Theorem} 
\newtheorem{definition}{Definition} 
\newtheorem{remark}{Remark}

\renewcommand{\vec}[1]{\boldsymbol{#1}}

\newcommand{\V}{\mathcal{V}}
\newcommand{\R}{\mathbb{R}}

\newcommand{\N}{\mathbb{N}}
\newcommand{\ReLU}{\mathrm{ReLU}}
\newcommand{\FF}{\mathsf{FF}}

\newcommand{\tf}{\mathsf{TF}}
\newcommand{\tm}{\mathsf{TM}}
\renewcommand{\pm}{\mathsf{PM}}
\newcommand{\emb}{\mathsf{emb}}
\newcommand{\pos}{\mathsf{pos}}
\newcommand{\out}{\mathsf{out}}
\newcommand{\dec}{\mathsf{dec}}
\newcommand{\hardmax}{\mathsf{hardmax}}
\newcommand{\softmax}{\mathsf{softmax}}
\newcommand{\SPACE}{\mathsf{SPACE}}
\newcommand{\pspace}{\mathsf{PSPACE}}
\newcommand{\poly}{\mathsf{poly}}
\newcommand{\window}{\mathsf{WINDOW}}

\title{Constant Bit-size Transformers Are Turing Complete}
\author{
Qian Li\thanks{Shenzhen International  Center For Industrial  And  Applied  Mathematics, Shenzhen Research Institute of Big Data. Email: \texttt{liqian.ict@gmail.com}}
\quad
Yuyi Wang\thanks{CRRC Zhuzhou Institute \& Tengen Intelligence Institute. Email: \texttt{yuyiwang920@gmail.com}}
}
\date{}
\begin{document}
\maketitle

\begin{abstract}
We prove that any Turing machine running on inputs of arbitrary length can be simulated by a constant bit-size transformer, as long as the context window is sufficiently long. This improves previous works, which require scaling up either the model's precision or the number of parameters on longer inputs. Furthermore, we prove that the complexity class $\SPACE[s(n)]$ exactly characterizes the expressive power of a constant bit-size transformer with a context window of length $s(n)$. Our approach relies on simulating Post machines, a Turing-complete computational model. Post machines can be modeled as automata equipped with a queue, exhibiting computational behaviors naturally aligned with those of transformers. The behavioral similarity between transformers and Post machines may offer new insights into the mechanisms underlying the reasoning abilities of transformers.
\end{abstract}

\section{Introduction}\label{sec:intro}
Transformer-based large language models (LLMs) with chain-of-thought (CoT) steps \citep{gpt4,team2023gemini,anthropic2024claude3,dubey2024llama,deepseek} have demonstrated exceptional capabilities across a wide range of complex reasoning tasks, such as mathematical problem solving and code generation \citep{hendrycks2021measuring,jimenez2023swe,rein2023gpqa}. These remarkable empirical successes raise a fundamental question: What enables transformers to support such general reasoning capabilities, and what are the limits of this capability? Understanding this question is not only of theoretical interest but also of practical significance for further enhancing the reasoning abilities of transformers.

Motivated by this fundamental question, a line of theoretical works \citep{perez,bhat,ashish2,ashish,tengyu,prompt,zubic,depth,pencil} have studied the reasoning abilities of transformers through the lens of expressiveness. A central result from these works is that transformers (using CoT steps) are Turing complete; that is, transformers have sufficient expressive power to simulate arbitrary Turing machines (TMs). However, in all these results, the bit-size of the transformers (i.e., the total number of bits representing the model) is not constant: either the precision bit \citep{perez,bhat,ashish,prompt} or the embedding dimension \citep{tengyu} must grow with the input length, meaning that larger models are required to handle longer inputs. This scaling requirement raises a fundamental barrier to approaching transformer-based AGI \citep{lecun,GoldblumFRW24}, particularly for lifelong learning agents that interact with an open environment and accumulate an ever-growing history (e.g., the AIXI agent \citep{uai}), where the input length grows unboundedly over time.

This motivates the following critical question:
\begin{center}
\textit{Problem 1: Is it necessary to continue scaling up the bit size of transformers to handle longer inputs?}
\end{center}

\begin{table}[t]
\centering
\caption{Comparing the existing Turing-completeness proofs in terms of required precision, embedding dimension, effective window size, and CoT length. We focus on the nontrivial case where $t(n),s(n)\geq n$. We remark that these constructions typically do not employ the vanilla architecture but instead introduce specific modifications; see Appendix \ref{app:mod} for a summary.}
\label{tab:turing-completeness}
\resizebox{\linewidth}{!}{
\begin{tabular}{lcccccc}
\toprule
\textbf{Source} & \textbf{Precision} & \textbf{Dimension} & \textbf{Window} & \textbf{CoT per TM-step}\\
\midrule
\cite{perez} & $O(\log t(n))$ & $O(1)$ & $n+t(n)$ & $1$ \\
\cite{bhat} & unbounded & $O(1)$ & $n+t(n)$ & $1$ \\
\cite{ashish} & $O(\log t(n))$ & $O(1)$ & $n+t(n)$ & $1$ \\
\cite{tengyu}$^\ast$ & $O(1)$ & $O(\log t(n))$ & $O(t(n)\log t(n))$ & $O(\log t(n))$ (amortized)  \\
\cite{prompt} & $O(\log t(n))$ & $O(1)$ & $O(t(n)\log t(n))$ & $O(\log t(n))$ (amortized)\\
\textbf{This work} & $O(1)$ & $O(1)$ & $s(n)$ & $s(n)$\\
\bottomrule
\end{tabular}
}
\vspace{2pt}
\raggedright \footnotesize $^{\ast}$ Any multi-tape TM running in time $t(n)$ can be simulated by a Boolean circuit of size $O(t(n)\log t(n))$ \citep{arora2009computational}.
\end{table}

\subsection{Our contribution} 
We answer Problem 1 in the negative. We prove that \emph{constant bit-size} transformers\footnote{Transformers are allowed to generate arbitrarily long  CoTs before producing the final output.} are Turing complete. Specifically, given any TM, there exists a transformer with fixed numerical precision and a fixed number of parameters that can simulate the TM on inputs of arbitrary length, provided that the transformer's context window is sufficiently long. In particular, by applying it to the universal Turing machine, which takes as input a description $\tm$ of a TM and an input string $x$ and outputs $\tm(x)$, we can conclude that a single, constant bit-size transformer can compute any computable function, as long as the description of a relevant TM is loaded in the prompt.

\begin{remark}
In fact, it is not even necessary to load the task-specific TM description into the prompt or to pre-inject it during pre-training. Instead, the transformer can be instantiated to simulate a TM that performs the meta-task of designing an algorithm for the given reasoning problem. For example, one can design the transformer to simulate Levin’s universal search algorithm \citep{levin1,levin2,oops}, which solves arbitrary search problems, such as theorem proving and planning, as quickly as the fastest possible algorithm in an asymptotic sense. This suggests that, in principle, a single constant bit-size transformer has sufficient expressive power to achieve a form of general reasoning ability.
\end{remark}

We emphasize that the context window length is the minimal resource needed to scale up to handle longer inputs: The context window must be at least as long as the input length $n$, otherwise the transformer cannot even access the entire input. Moreover, we provide an exact characterization of the expressive power of constant bit-size transformers as a function of their context window length. Specifically, let $\window[s(n)]$ denote the complexity class consisting of decision problems solvable by a constant bit-size transformer with a context window of length $O(s(n))$, $\SPACE[s(n)]$ the class of decision problems solvable by a TM using $O(s(n))$ space, and $\pspace:= \bigcup_{k \in \N} \SPACE[n^k]$ the class of decision problems solvable in polynomial space. We now state the main theorem of this paper:

\begin{restatable}{theorem}{maintheorem}\label{thm:main0}
For any non-decreasing function $s(n) \geq n$, we have $\window[s(n)] = \SPACE[s(n)]$. In particular, $\window[\poly(n)]=\pspace$.
\end{restatable}

Previous proofs establishing the Turing completeness of transformers suggested that the context window length must scale with the time complexity for solving the task. In contrast, Theorem \ref{thm:main0} implies that it is sufficient for the context window length to scale only with the space complexity instead. This distinction is crucial, since many problems exhibit a significant gap between space complexity and time complexity. For instance, solving  Boolean satisfication problems, Sokoban puzzles \citep{hearn2005pspace}, model checking tasks \citep{sistla1985complexity}, or other $\pspace$-complete problems, require only polynomial space by tracking the current configuration, whereas the time complexity is widely conjectured to be superpolynomial \citep{arora2009computational} or even $2^{\poly(n)}$ \citep{impagliazzo2001complexity}. Recently, \cite{ryan} proved that every multi-tape Turing machine running in time $t(n)$ can be simulated in space only $O(\sqrt{t(n)\log t(n)})$. As a consequence, there are problems solvable in $O(s(n))$ space that require $\tilde{\Omega}(s(n)^2)$ time on multitape Turing machines.


In addition, the simulation of a TM by a transformer in Theorem \ref{thm:main0} is uniform and easily generalizable with respect to the context window length. Specifically, to simulate a given TM on increasingly long inputs, the transformer just needs to extend its context window by adjusting the relative positional encoding according to an explicit formula, without changing all other model parameters.



\begin{remark}
Aiming for stronger deductive reasoning ability, a line of engineering efforts tend to lengthen context window instead of growing parameter count. Our result formally justifies this strategy: extending the window alone is sufficient. Moreover, a window that grows only polynomially with the input already suffices to solve every PSPACE problem, including game-tree search and mathematical proofs.
\end{remark}

\begin{remark}
There is a ongoing debta on whether transformers can approach AGI. A common negative view (e.g. \citep{lecun,GoldblumFRW24}) argues that transformers, limited by finite size and window, cannot approximate unbounded computation. Our result shows that, at least for deductive reasoning tasks (whose solutions depend only on explicit information), scaling window length alone is theoretically sufficient, and continually scaling up size is not a principled requirement.
\end{remark}

Our proof strategy is to simulate Post machines \citep{post1} with transformers. A Post machine can be modeled as an automaton equipped with a queue. By storing the TM's tape within the queue in a cyclic manner, Post machines can faithfully simulate Turing machines, and thus are Turing complete \citep{post1,davis1994computability}. By viewing the transformer's context window as a queue-like structure, one can observe that the behavior of Post machines naturally aligns with that of transformers. Leveraging this behavioral similarity, we can then design a transformer that faithfully simulates the step-by-step execution of a Post machine. In general, this alignment offers a new interpretation of the attention mechanism: not only as a statistical aggregator, but also as a discrete computational operation over queue-like structures.

\subsection{Other related work}
Recent theoretical studies have examined the computational capabilities of transformers under various architectural constraints.  
\cite{feng,ashish2,ashish,tengyu} proved that CoT steps can improve the reasoning capabilities of transformers: Transformers without CoT can only solve problems that can be solved by shallow circuits, whereas allowing polynomially long CoT steps enables the same model to solve any problem in $\mathrm{P}$. In addition, \cite{depth} showed that a transformer with depth growing only logarithmically in the input length can recognize regular languages and solve graph connectivity problems without CoT, whereas a constant-depth transformer requires CoT steps of length $\Omega(n)$. We remark that these papers assume either a $O(\log n)$-bit precision or a $O(\log n)$-embedding size, and require the context window to be long enough to cover the entire context.
The survey by \cite{strobl2024formal} provides a comprehensive overview on formal-language expressivity of transformers, clarifying how
design choices affect expressivity.

Aiming to reduce the required context window length and the memory, \cite{pencil} proposed the PENCIL framework, which incorporates a reduction mechanism to recursively ``erase'' intermediate CoT steps. This allows the context window length to scale with the space complexity of the computation, rather than the time complexity. We remark that their results still require increasing the bit-size of the transformer as the input length grows.

\cite{schuurmans2024autoregressive} considered a generalization of autoregressive decoding where, given a long input, emitted tokens are appended to the end of the sequence as the context window advances, and proved that the resulting system is Turing complete. This work also exploits the similarity between the behaviors of Post machines and Transformers. In fact, they analyze an even restrictive model, the so-called Lag system, where the next token depends solely on the leftmost \textbf{two} tokens in the queue and \textbf{not} on the state $q$ of PM. The main technical challenge in their construction is thus to compensate for the absence of access to the state $q$. As a result, their simulation requires $\Theta(s(n)^3)$ CoT tokens per TM step, whereas ours are only $O(s(n))$.



\section{Notations and preliminaries}\label{sec:notation}

Let $\R$ be the set of real numbers and $\N$ be the set of natural numbers. For $n\in \N$, let $[n]$ denote $\{1,2,\cdots,n\}$. We use bold lowercase letters to represent vectors or sequences (e.g., $\vec{x}$), and bold uppercase letters to represent matrices (e.g., $\vec{A}$). Given a vector $\vec{x}$ and indices $j \leq i$, we use $\vec{x}_{j:i}$ to denote the sub-vector $(x_j, x_{j+1}, \ldots, x_i)$. Let $\vec{0}_n$ denote the $n$-dimensional all-zeros vector. 
Given a vocabulary $\V$, let $\V^n$ denote the set of length-$n$ strings over $\V$, and let $\V^*:=\bigcup_{n=1}^{+\infty} \V^n$ denote the set of all finite strings over $\V$.

\subsection{Turing machines}

A (single-tape) Turing machine (TM) is defined as a tuple $\langle\Sigma,Q,\delta\rangle$ where 
\begin{itemize}
\item $\Sigma$ is the tape alphabet, including a designated ``blank'' symbol $\perp$, a designated ``start'' symbol $\triangleright$, and numbers 0 and 1.
\item $Q$ is a finite set consisting of possible states. We assume $Q$ contains a designated start state $q_{start}$ and a designated halting state $q_{halt}$.
\item $\delta: Q\times \Sigma\rightarrow Q\times \Sigma\times \{\mbox{Left},\mbox{Right}\}$ is a transition function describing the rules used in performing each step of the TM.
\end{itemize}
The machine has a single tape that is infinite to the right and bounded on the left. The tape is equipped with a tape head. Initially, the tape (from the leftmost cell rightward) contains the start symbol $\triangleright$, a finite $\{0,1\}$-string $x$ (the input), and the blank symbols $\perp$ on the rest of its cells. The head starts at the leftmost tape cell, and the machine starts in the state $q_{start}$. The computation then repeats the following step until the machine enters $q_{halt}$: if the machine is in some state $q\in Q$ and is reading symbol $\sigma$ from the tape, and if $\delta(q,\sigma)=(q',\sigma',z)$ for some $z\in\{\mbox{Left},\mbox{Right}\}$, then the machine replaces $\sigma$ with $\sigma'$ in the current cell, changes state from $q$ to $q'$, and moves the tape head one cell in direction $z$.

\paragraph{Space complexity}
Let $s:\mathbb{N}\to \mathbb{N}$ be a non-decreasing function. The complexity class $\SPACE[s(n)]$ consists of all decision problems $\{0,1\}^*\rightarrow\{0,1\}$ that can be decided by a single-tape Turing machine using at most $O(s(n))$ tape cells on inputs of length $n$.\footnote{We implicitly assume that $s(n)\geq n$, since $n$ tape cells are required to load an input of length $n$.} The class $\pspace:= \bigcup_{k \in \mathbb{N}} \mathsf{SPACE}[n^k]$ is defined as the class of decision problems that can be solved in polynomial space.

\subsection{Post machines}
A Post machine (PM) \citep{post1} is defined as a tuple $\langle\Sigma,Q,\delta\rangle$ where 
\begin{itemize}
\item $\Sigma$ is the tape alphabet, including a designated ``blank'' symbol $\perp$, a designated ``start'' symbol $\triangleright$, and numbers 0 and 1.
\item $Q$ is a finite set consisting of possible states. We assume $Q$ contains a designated start state $q_{start}$ and a designated halting state $q_{halt}$.
\item $\delta: Q \times \Gamma \to Q \times \Gamma\times \{\mbox{Stay},\mbox{Right}\}$ is the transition function, specifying how the machine updates the tape and state.
\end{itemize}
The machine has a single tape that is infinite to the right and bounded on the left. The tape is equipped with two tape heads: the \emph{front head} and the \emph{rear head}. Initially, the tape (from leftmost cell rightward) contains the start symbol $\triangleright$, a finite $\{0,1\}$-string $x$ (the input), and the blank symbols $\perp$ on the rest of its cells. The front head starts at the leftmost tape cell, the rear head starts at the right end of the input, and the machine starts in state $q_{start}$. The computation repeats the following step until the machine enters $q_{halt}$: if the machine is  in state $q\in Q$ with the front head reading symbol $\sigma$ from the tape, and if $\delta(q,\sigma)=(q',\sigma',z)$ for some $z\in\{\mbox{Stay},\mbox{Right}\}$, then the machine (i) changes the state from $q$ to $q'$, (ii) moves the front head either one cell to the right if $z=\mbox{Right}$, and (iii) moves the rear head one cell to the right, and then write $\sigma'$. 

One can see that a PM can also be modeled as an automaton equipped with a queue of unbounded size: the queue is initialized as the input $x$, the front head points to the first element of the queue, and the rear head points to the last one. In each step, the PM reads and deletes the first element of the queue and may append a string to the tail.

\paragraph{Turing completeness of PM} 
The model of PM is Turing complete, i.e., equivalent in computational power to TMs. Specifically, 
\begin{theorem}\label{thm:pmtm}
Given any single-tape TM running in $t(n)$ time and $s(n)$ space, there is an equivalent PM running in $O(t(n)\times s(n))$ time and using a $s(n)$-size queue.
\end{theorem}
This theorem has been implicitly proved in several places, e.g., \citep{pmtm2} and Section 9.1 of \citep{pettorossi2014elements}. For completeness, we present a proof in Appendix \ref{app:pmtm}.
The intuition is as follows. The PM's queue stores the non-blank part of the TM's tape in a cyclic manner: If the TM head changes a symbol $\sigma$ to $\sigma'$ and moves right, the queue shifts by deleting $\sigma$ from the front and appending $\sigma'$ to the end. If the TM head moves left, the queue cyclically rotates by moving the last symbol to the front, where the appended symbol may need to depend on the leftmost \textbf{two} queue symbols. 
To avoid explicitly pre-reading the next queue cell, we apply a one-step \emph{delayed append} trick: initially, pop the leftmost queue symbol and store it in the finite-state controller, without appending; thereafter, at each step, (i) pop the next leftmost symbol, and (ii) append a symbol determined by the two most recently popped symbols.

\subsection{Transformers}\label{sec:transformer}
For simplicity of notation, we present the definition of a single-head transformer, since only single-head attention is used in our proof. The definition of general multi-head transformers is deferred to Appendix~\ref{app:trans}.

Let $\mathcal{V}$ be a finite vocabulary. A decoder-only transformer is a parameterized neural network $\tf_{\theta}$ mapping from $\V^*$ to $\V$. 
Specifically, a transformer is a composition $\tf:=\out \circ \dec_{L-1} \circ\cdots\circ \dec_{0} \circ \pos \circ \emb$ of four kinds of layers. Given a sequence $\vec{v}=(v_1,\cdots,v_i)\in \V^i$, it computes the next token $v_{i+1}$ as follows:

\textbf{1. Token embedding layer ($\emb$)}: For $j\leq i$, map each $v_j\in \V$ to a vector $\emb(v_j)\in \R^d$. Here, $d$ is called the \emph{embedding size}.

\textbf{2. Positional encoding layer ($\pos$)}: For $j\leq i$, add a positional encoding $\pos(i-j)\in\R^d$ to the token embedding $\emb(v_j)$, resulting in the initial input representation $\vec{h}^{0}_j:=\emb(v_j)+\pos(i-j)$. Here, we assume a relative positional encoding scheme, where the encoding depends only on the distance between $v_j$ and $v_i$. 

\textbf{3. Decoder Layer ($\dec_\ell$ for $\ell=0,\cdots, L-1$)}: Each decoder layer $\dec_{\ell}$ consists of two sublayers: an self-attention layer, followed by a fully-connected feed-forward network. Residual connections 
are applied around each sub-layer \footnote{For simplicity, we ignore layer normalizations from the standard transformer architecture \citep{radford2019language} in our analysis. With a little more technical treatment as in \citep{tengyu}, our results can be extended to transformers with layer normalizations.}. Here, following common practice \citet{perez,ashish,ashish2,prompt,pencil}, we use $\hardmax$ as a realistic abstraction of $\softmax$. In fact, $\hardmax$ is the instance-wise limit of zero temperature limit of $\mathsf{softmax}$ \citep{pencil}. Consider a vector $\vec{s} \in \R^n$. If the maximum value in $\vec{s}$ appears at $t$ positions, then the $\hardmax$ function is defined coordinate-wise as:
\[
\hardmax(\vec{s})_j:= 
\begin{cases}
\frac{1}{t}, & \text{if } s_j = \max_k s_k, \\
0, & \text{otherwise}.
\end{cases}
\]


\textit{3.1. Single-head self-attention layer}: Compute the attention score
\[
\vec{s}_{i}^\ell=\hardmax\left(\langle \vec{h}^\ell_1\cdot \vec{Q}^{\ell},\vec{h}^\ell_i\cdot \vec{K}^{\ell} \rangle,\cdots,\langle \vec{h}^\ell_i\cdot \vec{Q}^{\ell},\vec{h}_i^\ell\cdot \vec{K}^{\ell} \rangle\right),
\]
and then
\[
\vec{a}^{\ell}_{i}=\sum_{j=1}^{i} s_{i,j}^\ell\cdot v^{\ell}(\vec{h}^{\ell}_i), \mbox{ where } v^{\ell} (h):=\vec{h}\cdot \vec{V}^{\ell} \mbox{ and } s^\ell_{i,j} \mbox{ is the } j\mbox{-th entry of } \vec{s}_{i}^\ell. 
\]
Here, $\vec{Q}^{\ell},\vec{K}^{\ell},\vec{V}^{\ell}\in\mathbb{R}^{d\times d}$ are parametrized matrices. Then, this sublayer returns 
\[
\vec{h}_i^{\ell+0.5}:=\vec{W}^\ell\cdot \vec{a}^{\ell}_i + \vec{b}^\ell + \vec{h}_{i}^{\ell},
\]
where $\vec{W}^\ell\in\R^{d\times d}$ and $\vec{b}^\ell\in \R^d$ are parametrized.

\textit{3.2. Feed-forward network}: 
Apply a multi-layer fully-connected $\ReLU$ neural network $\FF$ to $\vec{h}^{\ell+0.5}_i$, and returns
\[
\vec{h}^{\ell+1}_i=\FF(\vec{h}_i^{\ell+0.5})+\vec{h}_i^{\ell+0.5}.
\]

\textbf{4. Output layer ($\out$)}: The final output representations $\vec{h}^{L}_i$ are projected onto the vocabulary space using a linear transformation followed by a $\arg\max$ function:
\[
v_{i+1}:=\out(\vec{h}^L_i)=\arg\max(W^{\out}\cdot \vec{h}^L_i + \vec{b}^{\out}), \mbox{ where } W^{\out}\in \R^{|\V|\times d} \mbox{ and } \vec{b}^\out\in\R^{|\V|}.
\]


We say that a transformer operates with a \emph{context window} of length $s$ if the generation of $v_{i+1}$ depends only on the last $s$ tokens $v_{i-s+1},\cdots,v_i$. A transformer is said to be of $p$-bit precision if each parameter is represented using $p$ bits. We define the \emph{bit-size} of a transformer $\tf_{\theta}$ as $p \times |\theta|$, meaning that the entire model can be represented by using $p \times |\theta|$ bits. In this paper, when we refer to transformers, we mean transformers that are allowed to generate arbitrarily long immediate CoT steps, unless stated otherwise.

\begin{definition}\label{def:window}
We define $\window[s(n)]$ as complexity class consisting of the problems $\{0,1\}^*\rightarrow\{0,1\}$ that can be solved by a constant bit-size transformer with context window of length $O(s(n))$ on inputs of length $n$.
\end{definition}

\section{Proof of main results}\label{sec:proof}
In this section, we prove Theorem \ref{thm:main0}, which characterizes the expressive power of constant bit-size transformers with context window of length $s(n)$ in terms of the complexity class $\SPACE[s(n)]$. In particular, it implies that constant bit-size transformers are Turing complete. 


First, we prove that any constant bit-size transformer with a context window of length $s(n)$ can be simulated by a Turing machine using $O(s(n))$ space.
\begin{theorem}\label{thm:upper}
For any non-decreasing function $s(n)\geq n$, $\window[s(n)]\subseteq \SPACE[s(n)]$.
\end{theorem}
\begin{proof}
Let $L:\{0,1\}^*\to\{0,1\}$ be a decision problem that can be solved by a constant bit-size transformer using a context window of length $s(n)$. We sketch a Turing machine using $O(s(n))$ space that solves $L$ by simulating the transformer. The Turing machine maintains a buffer of size $s(n)$ storing the latest $s(n)$ tokens in the transformer's context. To compute the next token generated by the transformer, the Turing machine simulates the computation procedure described in Section \ref{sec:transformer}, using an additional $O(s(n))$ working space. After generating the next token, the Turing machine updates the buffer by appending the new token and deleting the oldest token that slides out of the window, and then cleans the working space. By iteratively repeating this procedure, one can see that the Turing machine can faithfully simulate the transformer's behavior and solve the problem $L$.
\end{proof}

In the following, we prove the reverse direction, which is the main technical part.
\begin{theorem}\label{thm:lower}
Let $\tm$ be a single-tape Turing machine that, on input $x\in\{0,1\}^n$, uses at most $s(n)$ space and runs for at most $t(n)$ steps. There exists a constant bit-size transformer with a context window of length $O(s(n))$ that, on input $x$, takes $O(t(n)\cdot s(n))$ CoT steps and then outputs $\tm(x)$.

As a consequence, we have $\window[(s(n))]\supseteq \SPACE[s(n)]$ for $s(n)\geq n$.
\end{theorem}
\begin{proof}
Let $\tm$ be a single-tape Turing machine that runs in $t(n)$ steps and uses at most $s(n)$ tape cells. By Theorem \ref{thm:pmtm}, the $\tm$ can be simulated by a Post machine 
that runs in $O(t(n) \cdot s(n))$ time and uses a queue of size $s(n)$. We slightly adapt the Post machine to ensure that the queue size remains exactly $s(n)$ throughout the computation, or equivalently, that the distance between the front and rear heads on the tape remains fixed at $s(n)$. Specifically, we pad the input $x$ with $s(n)-n$ copies of a special symbol $\#$ to the right initially, and in each computation step, both the front and rear heads always move one cell to the right. 
One can easily see that such an adapted $\pm = (\Sigma,Q,\delta)$ simulates the Turing machine as well. Furthermore, without loss of generality, we assume $\Sigma = \{0,1,\#\}$, $Q=\{0,1\}^c$ for some $c\in\N$, and that $q_{start}\notin \delta(\sigma,q)$ for any $(\sigma,q)$, meaning that once the computation starts, the Post machine leaves $q_{start}$ immediately and never comes back.   


We now construct a constant bit-size transformer $\tf$ with a context window of length $s(n)$ to faithfully simulate $\pm$. The intuition is as follows:
\begin{itemize}
\item The $s(n)$-size queue is simulated by the $s(n)$-long context window, where the first element in the queue (or the tape cell pointed by the front head) 
corresponds to the oldest token in the window, and the last element in the queue (or the tape cell pointed by the rear head) corresponds to the newest token. 
Besides, we will let the vocabulary of the transformer be $\V=\Sigma\times Q$, so that a token also tracks the $\pm$ state information. 
\item As mentioned above, in each computation step of $\pm$, the first element in the queue is removed, and a new element is added, so that the queue size remains the same. Correspondingly, in each CoT step of the transformer, the oldest token in the window slides out, and a new token is appended.
\item The transition function of $\pm$ takes the last element in the queue and the current state as input, and outputs the next state and the next element. Correspondingly, by carefully choosing the parameters, the self-attention layer retrieves the last element from the oldest token in the window, and the current state from the current token; subsequently, a feed-forward network is used to implement the transition function $\delta$. 
\end{itemize}

We formally define the transformer $\tf$ as follows: let $\V=\Sigma\times Q=\{0,1,\#\}\times \{0,1\}^c$. 

\underline{1. Token embedding layer}\quad Map each token $v=(\sigma,q)$ to a vector $\emb(\sigma,q)\in \R^{c+8}$ as follows: 
\begin{flalign}\label{eq:emb}
\emb(\sigma,q)=\begin{cases}
 (1,1,0,0,q,0,0,0,0), & \text{if } \sigma=0;\\
 (1,0,1,0,q,0,0,0,0), & \text{if } \sigma=1;\\
 (1,0,0,1,q,0,0,0,0), & \text{if } \sigma=\#.
 \end{cases}
 \end{flalign}


\underline{2. Positional encoding layer}\quad Add a relative positional encoding $\pos(i-j)\in\R^{c+8}$ to the token embedding $\emb(v_j)$ where
\begin{flalign}\label{eq:pos}
\pos(i-j)=\begin{cases}
 \vec{0}_{c+8}, & \text{if } 0<i-j<s(n)-1;\\
 (\vec{0}_{c+7},1), & \text{if } i-j=0,\\
 (\vec{0}_{c+7},2), & \text{if } i-j=s(n)-1,
 \end{cases}
 \end{flalign}
Then return $\vec{h}_j^0=\emb(v_j)+\pos(i-j)$.

\underline{3. Decoder layer}\quad The transformer has only one decoder layer, and the decoder layer has only one attention head. In the self-attention layer, we set the matrix $\vec{K},\vec{Q},\vec{V},\vec{W}\in\R^{d\times d}$ and $\vec{b}\in\R^d$ where $d=c+8$ as follows:
\begin{itemize}
\item Matrix $\vec{K}$: it has only one non-zero entry $K_{c+8,1}=1$;
\item Matrix $\vec{Q}$: it has only one non-zero entry $Q_{c+8,c+8}=1$;
\item Matrix $\vec{V}$: it has four non-zero entries: $V_{c+5,2}=V_{c+6,3}=V_{c+7,4}=V_{c+8,c+8}=1$.
\end{itemize}
Besides, we set the matrix $\vec{W}\in \R^{d\times d}$ to be the identity matrix, and $\vec{b}=\vec{0}_{c+8}$.


The feed-forward network $\FF$ is designed to 
simulate transition function $\delta$. Specifically, we let
\begin{flalign}\label{eq:ff}
\FF(\vec{h})+\vec{h}=\begin{cases}
 \vec{e}_{(\#,q_{start})}, & \text{if } h_{c+8}=2;\\
 \vec{e}_{\delta(0,q)}, & \text{if } \vec{h}_{4:c+8}=(q,1,0,0,3);\\
 \vec{e}_{\delta(1,q)}, & \text{if } \vec{h}_{4:c+8}=(q,0,1,0,3);\\
 \vec{e}_{\delta(\#,q)}, & \text{if } \vec{h}_{4:c+8}=(q,0,0,1,3);
\end{cases}
\end{flalign}
Here, $\vec{e}_{(\sigma',q')}\in\R^{|\V|}$ is the unit vector where the entry indexed by $(\sigma',q')$ is 1 and any other entry is 0.


\underline{4. Output layer}\quad We set $W^{\out}$ to be the identity matrix and $\vec{b}^{\out}$ the all-zeros vector. Then the output layer outputs
$v_{i+1}:=\arg\max(\vec{h}^1_i)$. 

In the following, we show that the transformer $\tf$ faithfully simulates the Post machine $\pm$. Let $(\sigma_1,q_1),(\sigma_2,q_2),\cdots$ denote the execution log of $\pm$ running on $x\in\{0,1\}^n$, where $\sigma_i$ is the $i$-th element added to the queue and $q_i$ is the state when adding $\sigma_i$. Note that $(\sigma_i,q_i)=(x_i,q_{start})$ for $i\leq n$, $(\sigma_i,q_i)=(\#,q_{start})$ for $n+1\leq i< s(n)$, and $\delta(\sigma_{i-s(n)+1},q_i)=(\sigma_{i+1},q_{i+1})$ for $i\geq s(n)$. 
Let $v_1, v_2, \dots$ denote the token sequence in the context of $\tf$ when it takes $v_1 = (x_1, q_{start}), \dots, v_n = (x_n, q_{start})$ as input.
 We show by induction that for each $i\geq 1$, we have $v_i=(\sigma_i,q_i)$.

\underline{Base case: $i\leq n$}\quad This case is trivial. 

\underline{Inductive case: $n+1\leq i<s(n)$}\quad  In the self-attention layer, observing that $\vec{K}\cdot \vec{h}^0_i=(\vec{0}_{c+7},1)=1$ and 
\begin{flalign}
\vec{Q}\cdot \vec{h}^0_j=\pos(i-j)=\begin{cases}
 \vec{0}_{c+8}, & \text{if } 1\leq j<i,\\
 (\vec{0}_{c+7},1), & \text{if } j=i,
 \end{cases}
\end{flalign}
we have
\begin{align*}
\vec{s}_{i} = &\hardmax\left(\langle \vec{h}^0_{1}\cdot \vec{Q},\vec{h}^0_i\cdot \vec{K}\rangle,\cdots,\langle \vec{h}^0_i\cdot \vec{Q},\vec{h}^0_i\cdot \vec{K}\rangle\right)=(0,\cdots,0,1)\in\R^i
\end{align*}
and
\[
\vec{a}_{i}=\sum_{j=1}^{i} s_{i,j}\cdot \left(\vec{V}\cdot \vec{h}^0_j\right)=\sum_{j=1}^{i} s_{i,j}\cdot \left(\vec{0}_{c+4},\vec{h}^0_{j,2;4},h^0_{j:c+8}\right)=(\vec{0}_{c+4},\vec{h}_{i,2:4}^0,1).
\]
Then 
\[
\vec{h}_i^{0.5}=\vec{W}\cdot \vec{a}_i+\vec{b}+\vec{h}_i^0=\vec{a}_i+\vec{h}_i^0=(\vec{h}^0_{i,1:c+4},\vec{h}^0_{i,2:4},2).
\]

In the feed-forward network layer, $\vec{h}_i^{0.5}$ is mapped to $\FF(\vec{h}_i^{0.5})+\vec{h}_i^{0.5}$, which is $\vec{e}_{(\#,q_{start})}$ according to Equation (\ref{eq:ff}). Finally, the output layer outputs $(\#,q_{start})$ as desired.

\underline{Inductive case: $i\geq s(n)$}\quad  In the self-attention layer, observing that $\vec{K}\cdot \vec{h}^0_i=(\vec{0}_{c+7},1)=1$ and 
\begin{flalign}
\vec{Q}\cdot \vec{h}^0_j=\pos(i-j)=\begin{cases}
 0, & \text{if } i-s(n)+1<j<i;\\
 1, & \text{if } j=i,\\
 2, & \text{if } j=i-s(n)+1,
 \end{cases}
\end{flalign}
we have
\begin{align*}
\vec{s}_{i} =\hardmax\left(\langle \vec{h}^0_{i-s(n)+1}\cdot \vec{Q},\vec{h}^0_i\cdot \vec{K}\rangle,\cdots,\langle \vec{h}^0_i\cdot \vec{Q},\vec{h}^0_i\cdot \vec{K}\rangle\right)=(1,0\cdots,0)\in \R^{s(n)}
\end{align*}
and
\[
\vec{a}_{i}=\sum_{j=i-s(n)+1}^{i} s_{i,j}\cdot \left(\vec{0}_{c+4},\vec{h}^0_{j,2;4},h^0_{j:c+8}\right)=(\vec{0}_{c+4},\vec{h}_{i-s(n)+1,2:4}^0,2).\]
Then 
\[
\vec{h}_i^{0.5}=\vec{a}_i+\vec{h}_i^0=(\vec{h}^0_{i,1:c+4},\vec{h}^0_{i-s(n)+1,2:4},3).
\]

In the feed-forward network layer, noting that by the induction hypothesis, we have
\[
\vec{h}_{i,5:c+4}^{0.5}=\vec{h}_{i,5:c+4}^{0}=q_{i}, \text{ and } \vec{h}_{i,c+5:c+7}^{0.5}=\vec{h}^0_{i-s(n)+1,2:4}=\sigma_{i-s(n)+1}.
\]
The vector $\vec{h}_i^{0.5}$ is mapped to $\FF(\vec{h}_i^{0.5})+\vec{h}_i^{0.5}$, which is $\vec{e}_{\delta(\sigma_{i-s(n)+1},q_i)}$ according to Equation (\ref{eq:ff}). Recall that $\delta(\sigma_{i-s(n)+1},q_i)=(\sigma_{i+1},q_{i+1})$, then one can see that the output layer outputs $(\sigma_{i+1},q_{i+1})$ as desired. 

Now, we have shown that $v_i=(\sigma_i,q_i)$ for any $i\geq 1$ and thus finished the proof.
\end{proof}

\begin{remark}
As we can see from the proof, to simulate the given TM on longer inputs, what we need to do is just change the relative positional encoding according to Equation (\ref{eq:pos}), without changing all other model parameters.
\end{remark}
\begin{remark} We briefly compare our construction with existing Turing-completeness proofs and highlight why our construction achieves constant bit-size. In \citep{perez} and related follow-ups, the query may attend to many earlier tokens; when multiple positions tie for the maximum score, the hardmax outputs fractions such as $1/t$, which requires the precision to grow with $t$. The proof of \cite{tengyu} encodes the simulated circuit directly into the absolute positional encoding, rather than into the parameters. Consequently, as the circuit expands, the positional encoding must also grow in dimension, so the embedding size is no longer constant. 

In contrast, our construction leverages the automaton-like behavior of Post machines, which is simpler and more regular than the head movements of Turing machines or the wiring patterns of circuits. As a result, at each CoT step the query attends to exactly one token, specifically the token located $s(n)$ positions earlier. This fixed-offset attention can be implemented purely through a relative positional encoding, without any additional arithmetic operations.

In particular, the input to $\hardmax$ is always one-hot, so is the resulted attention distribution. As an alternative of $\hardmax$, one could replace $\hardmax$ with the right-most hardmax function \citep{DBLP:conf/nips/Yang0A24}.
\end{remark}

Combining Theorems \ref{thm:upper} and \ref{thm:lower}, we immediately conclude our main theorem. 
\maintheorem*


\section{Conclusions and Discussions}\label{sec:conclusion}
In this work, we have shown that a constant bit-size transformer, with a sufficiently long context window, can simulate any Turing machine on arbitrarily long inputs. We further proved that the complexity class $\SPACE[s(n)]$ precisely characterizes the expressive power of constant bit-size transformers with a context window of length $s(n)$, suggesting that the context window length needs to scale only with the space complexity rather than the time complexity as suggested by previous results. Our proof leverages the behavioral similarity between transformers and Post machine, which offers new insights into the mechanism underlying the reasoning abilities of transformers. We list some future directions:
\begin{itemize}[leftmargin=16pt]
\item \textit{Simulation efficiency}: Our construction requires $O(t(n)\cdot s(n))$ CoT steps, which is potentially prohibitive for practical applications. It would be interesting and important to investigate whether this slowdown can be avoided without compromising optimality in other aspects.
\item \textit{Positional encodings}: Our construction employs an nonstandard relative positional encoding. It is open whether it can be replaced with fixed absolute positional encodings or other standard relative positional encodings. Moreover, our positional encoding explicitly depends on the assumed space upper bound, and it is unclear how this bound could be inferred automatically.
\item \textit{Learnability and empirical validation}: Our contribution is purely about expressiveness. Whether standard training procedures can learn such behavior is an important open problem. How well our construction behaves in practice also remains to be investigated.
\end{itemize}

\bibliographystyle{abbrvnat}
\bibliography{main.bib}

\appendix
\section{Modifications on Transformers in Turing Completeness Proofs}\label{app:mod}

Existing proofs of Turing-completeness for Transformers typically departure from the vanilla architecture and make specific modifications. \citet{perez} consider a Transformer with an absolute positional encoding given by the triplet $(i,1/i,1/i^2)$ at position $i$. In their construction, the attention score is defined as the negative absolute value of the dot product, and the attention mechanism uses average-hard attention. Moreover, the feed-forward layers employ sigmoid activations in place of ReLUs. \citet{ashish} dispense with positional encodings altogether and instead adopt \emph{Strict Causal Masking}, where attention at position $i$ can access all tokens up to position $i-1$ but not the current token $i$. Their construction also uses average-hard attention and introduces a \emph{Projected Pre-Norm}: an arbitrary linear projection is allowed before layer normalization. In the proof of \citet{tengyu}, the simulated circuit is directly encoded into the absolute positional encoding, rather than into the parameters. \citet{prompt} consider Transformers with nonstandard absolute positional encodings. In their construction, the query, key, and value maps in the attention sublayer are implemented as ReLU networks rather than linear transformations. This work employ nonstandard \emph{relative} positional encodings.

\section{Proof of Theorem \ref{thm:pmtm}}\label{app:pmtm}
Let $\tm=(\Sigma,Q,\delta)$ be a single-tape Turing machine that runs in time $t(n)$ and space $s(n)$. We construct an equivalent Post machine $\pm=(\Sigma',Q',\delta')$ that simulates $\tm$ within $O(t(n)\cdot s(n))$ time and uses a queue of size at most $s(n)$. 

\paragraph{Queue alphabet} We define the queue alphabet $\Sigma':=\{\sigma,\hat{\sigma},\tilde{\sigma}\mid \sigma\in \Sigma\}$. Here, $\hat{\cdot}$ and $\tilde{\cdot}$ are temporary markers used for simulating a left move of $\tm$'s head. Specifically, $\hat{\cdot}$ temporarily marks the current head cell, while $\tilde{\cdot}$ temporarily marks its left neighbor.

\paragraph{Delayed append trick} $\pm$ maintains a \emph{logical queue} $L$ which represents the non-blank segment of $\tm$'s tape in a cyclic manner. Formally, we realize $L$ as $L=R\circ Q_{\text{phys}}$ where
\begin{itemize}[leftmargin=16pt]
  \item $R$ is a register in the finite control storing the leftmost symbol of $L$, and
  \item $Q_{\text{phys}}$ is the actual queue content. Thus $|Q_{\text{phys}}|=|L|-1$.
\end{itemize} 
With this representation, the appended symbol is no longer restricted to depend only on the leftmost symbol of $L$; it can now be determined by the two leftmost symbols.


\paragraph{Initialization} Initially, $\tm$'s tape contains the start symbol $\triangleright$, followed by the input $x$, and then blanks. Accordingly, the logical queue $L$ is initialized to $(\triangleright,x)$: the first symbol $\triangleright$ is placed in the register $R$, and the remaining symbols $x$ are stored into $Q_{\text{phys}}$. During the simulation, we will keep the following invariant:
\begin{itemize}[leftmargin=16pt]
\item Right before each simulated step, the leftmost cell of $L$ correponds to $\tm$'s head cell; its symbol is either unmarked or carries the $\tilde{\cdot}$ mark, while all other cells of $L$ are unmarked.
\end{itemize}

\paragraph{Simulation of one step} Suppose $\tm$ is in state $q$ and its head reads $\sigma$ from cell $i$. Let $\delta(q,\sigma)=(q',\sigma',z)$ with $z\in\{\mbox{Left},\mbox{Right}\}$. This means that $\tm$ overwrites cell $i$ with $\sigma'$, change state from $q$ to $q'$, and moves its head one cell in direction $z$.
\begin{itemize}[leftmargin=16pt]
\item \textit{Move Right}. $\pm$ deletes the leftmost logical symbol and appends $\sigma'$ to $L$, thereby overwriting tape cell $i$. If the resulting leftmost logical symbol is $\triangleright$, indicating that $\tm$ has extended its non-blank segment by one blank cell to the right, then $\pm$ additionally appends a blank symbol $\bot$.
\item \textit{Move Left}. $\pm$ deletes the leftmost logical symbol and appends $\hat{\sigma'}$ to $L$, thereby overwriting tape cell $i$. It then cyclically rotates $L$ until the second leftmost logical symbol carries the $\hat{\cdot}$ mark. At this point, the first and second of $L$ correspond to tape cells $i-1$ and $i$, respectively. $\pm$ then adds a mark $\tilde{\cdot}$ to the first symbol and removes the $\hat{\cdot}$ mark from the second, after one additionally cyclic sweep of $L$; concretely, 
\begin{enumerate}
\item delete the leftmost element and append the same symbol with the mark $\tilde{\cdot}$; then
\item delete the leftmost element and append the same symbol with no mark; then
\item cyclically rotate $L$ until the leftmost logical symbol carries the $\tilde{\cdot}$ mark.
\end{enumerate}
\end{itemize}

\paragraph{Analysis} By induction on the number of simulated steps, one can check that the invariant is maintained throughout the simulation, and hence $\pm$ faithfully simulate $\tm$. For a right move, $\pm$ performs $O(1)$ queue operations. For a left move, it performs two full sweeps of $L$, costing $O(|L|)=O(s(n))$ operations. Therefore, the overall running time is $O(t(n)\cdot s(n))$, completing the proof.

\section{Multi-head Transformers}\label{app:trans}
A multi-head Transformer employs the following \emph{multi-head self-attention layer} in place of the single-head self-attention layer, while all other components remain the same as in the single-head transformer.

\textit{Multi-head self-attention layer}: For each head $k=1,2,\cdots,H$, compute the attention score
\[
\vec{s}_{k,i}^\ell=\hardmax\left(\langle \vec{h}^\ell_1\cdot \vec{Q}_{k}^{\ell},\vec{h}^\ell_i\cdot \vec{K}_k^{\ell} \rangle,\cdots,\langle \vec{h}^\ell_i\cdot \vec{Q}_{k}^{\ell},\vec{h}_i^\ell\cdot \vec{K}_k^{\ell} \rangle\right),
\]
and then
\[
\vec{a}^{\ell}_{i,k}=\sum_{j=1}^{i} s_{k,i,j}^\ell\cdot v^{\ell}_k(\vec{h}^{\ell}_i), \mbox{ where } v^{\ell}_k (h):=\vec{h}\cdot \vec{V}^{\ell}_k
\]
Here, $\vec{Q}^{\ell}_k,\vec{K}^{\ell}_k,\vec{V}^{\ell}_k\in\mathbb{R}^{d\times d/H}$ are parametrized matrices. By concatenating the $H$ heads, we obtain $\vec{a}^{\ell}_i=\left((\vec{a}^\ell_{i,1})^T,\cdots,(\vec{a}^\ell_{i,H})^T\right)^T\in \R^d$. Then, this sublayer returns 
\[
\vec{h}_i^{\ell+0.5}:=\vec{W}^\ell\cdot \vec{a}^{\ell}_i + \vec{b}^\ell + \vec{h}_{i}^{\ell},
\]
where $\vec{W}^\ell\in\R^{d\times d}$ and $\vec{b}^\ell\in \R^d$ are parametrized.

\end{document}